Bricks have a thickness of 12cm. If an external wall of a house measures 10m x 15m, calculate the rate of heat exchange through this wall if the inside temperature is 20 degrees celsius and the outside temperature is 5 degrees celsius
1
Expert's answer
2012-03-16T09:25:49-0400
Use the formula: Q=atS(T1-T2)/H (a=0.5 - coefficient, t- time, T1-T2-temperature
difference, S- area, H-thickness) =>Q/t=as(T1-T2)/H=0.5*10*15*15/0.12=93.75